# Learning Diffeomorphism for Image Registration with Time-Continuous Networks using Semigroup Regularization

### v2

@@ 5 Experiments
5.1 Datasets and Preprocessing
To evaluate SGDIR and compare its performance with other methods, we consider four publicly available datasets on 3D MR brain and lung scans widely used in the literature, namely OASIS [43], CANDI [44], LPBA40 [45], and DIRLAB [46]. OASIS dataset contains 416 T1 weighted scans from subjects aging from 18 to 96 with 100 of them diagnosed with mild to moderate Alzheimer’s disease. The segmentation masks of 35 subcortical regions available in the dataset serve as the ground truth for further evaluation of the registration. CANDI dataset contains T1 weighted brain scans of 117 subjects divided into 4 different subgroups including Healthy Control (HC), Schizophrenia Spectrum (SS), Bipolar Disorder with Psychosis (BPDwithPsy), and Bipolar Disorder without Psychosis (BPDwithoutPsy). The segmentation masks of 32 subcortical regions available in the dataset are used as the ground truth for the evaluation. LPBA40 dataset comes with 40 scans and the segmentation maps of 56 cortical and subcortical regions. For these datasets we use the skull stripped, MNI152 1mm normalized images and center-crop the images to the size $160\times 192\times 144$. For all datasets we use all the available anatomical structures for evaluation. DIRLAB dataset contains 10 pairs of lung CT scans corresponding to 10 patients. Each pair consists of the scans of lungs in the extreme inhaling and exhaling phases. The dataset also comes with 300 landmark points per image, annotated manually by experts, which are used for the evaluation.
We use min-max intensity normalization to transform the voxel intensities into the $[0,1]$ range. In the training phase, in order for the arguments of Proposition 1 remain valid we add a slight standard Gaussian noise with 0 mean and 0.01 standard deviation to the data. This makes sure that no pairs of voxels have precisely the same intensity levels while the image pair remaining visually intact.
@@ 5.2 Experimental Settings
In the OASIS dataset we use the first 256 subjects for training the model, the next 50 subjects for the validation and the rest of the subjects for the test set. By random sampling we generate 900 pairs for training, 100 pairs for validation and 1000 pairs for test. In the CANDI dataset, we mix all the subgroups and then use 80 subjects for training, 11 subjects for validation, and the rest for the test set. In a similar way to the OASIS dataset, we generate 400 pairs for training, 25 pairs for validation, and 300 pairs for test. For the LPBA40 dataset 190 pairs from 20 subjects are used as the training set, 10 pairs from 5 subjects are used as the validation set and 105 pairs from 15 subjects are used for the test set. For the DIRLAB dataset, 8 subjects are used for training and 2 subjects are used for test. We compare the performance of our model in terms of Dice score (for the OASIS, CANDI, and LPBA40 datasets), TRE (for the DIRLAB dataset), and the percentage of negative determinants of Jacobian of the deformation with a traditional optimization-based method, SyN [3], as one of the most successful methods for diffeomorphic image registration, learning-based methods including SYMNet [19], VoxelMorph [20], LapIRN [37], and DiffuseMorph [27], and optimization-based methods such as IDIR [7], NODEO [8], DNVF and its variant Cas-DNVF [9]. For SyN, we used the implementation in the DIPY package [47] and used cross-correlation for the loss function following the convention of the VoxelMorph. For other methods, except for DNVF/Cas-DNVF whose implementation was not available at the time of writing this paper, we used the codes available in their corresponding official repositories. The architecture of the UNet along with the time embedding modules is shown in Figure 1. The UNet encoder has 4 down-sampling layers of dimensions 32, 64, 128, 128, and 256, and the decoder has 4 up-sampling layers with the same dimensions as the down-sampling layers but in the reversed order. The time-embedding dimension is set to 64. All the activation functions for the layers are set to SiLU [48] to provide more smoothness to the network. In our experiments the regularization factor for the semigroup term is set to $\lambda=5e^{5}$, the NCC window size is set to 11, and the training is carried out for 100 epochs using Adam optimizer with the learning rate $lr=1e^{-4}$. The whole implementation is done in PyTorch and is tested on both NVIDIA GTX 1080 Ti and NVIDIA Tesla V100 GPUs with 32Gb RAM.
5.3 Evaluation Metrics
Following the convention of evaluating diffemorphic image registration methods ([9, 8, 19, 37]), we use Dice Similarity Coefficient to measure the similarity of the registered and target segmentation maps for the brain scan dataset, and we use TRE metric to measure the error between the registered and target landmark points for the DIRLAb dataset. We also use the negative Jacobian determinant percentage to measure the amount of disturbance in the topology of the moving image.
To compute the Dice score, we follow the convention of VoxelMorph [20] for computing the Dice score, which measures the mean Dice over all anatomical structures (35, 32, and 56 anotimical structures for the OASIS, CANDI, and LPBA40 datasets, respectively). For the DIRLAB dataset, the images are resampled to isotropic voxels of size $1mm$, and the TREs are reported in terms of average error in millimeter.
@@ 5.4 Results
Tables 1 and 2 provide a comprehensive comparison of the performance of SGDIR with other aforementioned methods on OASIS, CANDI, LPBA40, and DIRLAB datasets, respectively. Table 1 provides the performance of models in terms of Dice score and the percentage of negative Jacobian determinants. All the values are expressed as their average over the test set along with their standard deviation. For DNVF and Cas-DNVF due to unavailability of the implementation the results are directly taken from the corresponding paper. We consider the percentage of negative Jacobian determinants smaller than $10^{-5}\%$ negligible and show it as $\approx 0$ in the tables. Also, Figure 2 illustrates the qualitative comparison between our model and several other models on the OASIS dataset in terms of the warped image, warped segmentations, and the warping grids. Figure 3 illustrates the visual comparison between SGDIR and IDIR over the DIRLAB dataset.
Table 1 indicates that SGDIR outperforms competent methods both in terms of Dice score, and $|det(J_{\phi})<0|$, and Table 2 demonstrates the advantage of SGDIR in terms of the TRE metric. This implies that the imposition of the semigroup property as a regularization in the time continuum of the deformation field can in fact lead to a smooth deformation.
Among the compared methods, DiffuseMorph is also capable of generating a temporal deformation flow. Therefore, we have provided a visual comparison of warped images within 7 time steps for the OASIS dataset in Figure 4. In this example we have used the subject id 1 and 2 of the OASIS dataset as the fixed image and the moving image respectively. More visual results are also provided in Appendix 7.3.
@@
Table 1: Comparison of the performance of SGDIR with other methods on the OASIS, CANDI, and LPBA40 datasets. The values inside the parenthesis indicate the standard deviation. Cells with - indicate that the corresponding information is missing in the literature. $\approx 0$ sign is used for negative determinant percentages smaller than an order of $10^{-5}$.
OASIS CANDI LPBA40
Category Model Dice ($\uparrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$) Dice ($\uparrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$) Dice ($\uparrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 78.14 (1.32) 0.0531% (0.0241) 73.60 (1.15) 0.0011% (0.0008) 66.82 (3.17) 0.0194% (0.0072)
Learning VoxelMorph 75.93 (0.044) 0.0026% (0.0027) 75.63 (1.78) 0.0011% (0.0010) 70.72 (2.33) 0.0023% (0.0014)
SYMNet 79.39 (1.74) 0.0043% (0.0017) 75.89 (1.13) 0.0015% (0.0013) 69.98 (1.37) 0.0018% (0.0012)
LapIRN 78.80 (1.92) 0.0011% (0.0003) 76.23 (1.76) 0.0004% (0.0003) 71.88 (1.36) $\mathbf{\approx 0}$
DiffuseMorph 76.23 (2.23) 0.0551% (0.0431) 75.63 (2.01) 0.0041% (0.0012) 68.73 (2.88) 0.0032% (0.0011)
Optimization IDIR 72.31 (1.239) 0.0389% (0.0025) 74.42 (1.12) 0.0113% (0.0081) 67.93 (2.03) 0.0013% (0.0008)
NODEO 80.86 (1.14) 0.0105% (0.0115) 78.02 (1.52) $\approx$ 0% 72.31 (1.12) 0.0009% (0.0004)
DNVF 79.4 0.0015% 78.6 0.0003% - -
Cas-DNVF 81.5 0.0036% - - - -
Our Method SGDIR 82.23 (1.3851) 0.0009% (0.0007) 80.73 (1.25) $\mathbf{\approx 0}$74.23 (2.15) 0.0007% (0.0003)
Figure 2: Qualitative comparison of SGDIR with other methods on the OASIS dataset in terms of the warped images (first row), warped segmentations (second row) and warping grids (third row).
@@
Table 2: Comparison of the performance of SGDIR with other methods on the DIRLAB dataset. TRE is reported in terms of millimeters which is the same as the voxel size
Category Model TRE (mm) ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 1.35 0.0034%
Learning VoxelMorph 1.33 0.0011%
SYMNet 1.27 0.0006%
LapIRN 1.21 0.0001%
DiffuseMorph 1.35 0.0010%
Optimization IDIR 1.13 0.0005%
NODEO 1.08 0.0003%
Our Method SGDIR 0.98 0.0003%
Figure 3: Visual comparison of SGDIR with IDIR over the DIRLAB dataset. Deformation fields are depicted over the landmark points inside the lungs.
@@
Figure 4: The visual comparison of the evolution of the deformation between DiffuseMorph and SGDIR on the OASIS dataset over 7 time steps.
@@ 5.5 Ablation Study
In addition to the main results on the efficacy of SGDIR in providing high Dice scores while preserving the topology, we investigate the effect of some hyperparameters and design choices along with some consequent properties of SGDIR. More precisely, we investigate the effect of the semigroup regularization factor $\lambda$, the effect of continuous time scale against discrete time, and the topology-preserving property throughout the time interval. Since the behaviour of the model is the same for all datasets, we provide the analysis only for the OASIS dataset which is a more complex dataset with larger deformations.
@@ 5.5.1 Effect of Regularization Factor
Eq. (13) is responsible for imposing the semigroup property to the model and consequently, make the model produce topology-preserving non-folding deformations. Table 3 provides the effect of changing the factor $\lambda$ for the semigroup regularization on the Dice score and $|det(J_{\phi})<0|$. This table indicates that for a choice of a very large $\lambda$ we could achieve zero-folding deformations with a decrease in the dice score. Decreasing $\lambda$ gives more freedom to the model to increase the similarity of the deformed image and the fixed image in exchange for a larger folding percentage. An interesting property exhibited here is that drastically decreasing $\lambda$ (e.g., $\lambda=1$) is not helpful for achieving the highest Dice score. This indicates that the semigroup property is also helpful in narrowing down the trajectory space of the deformations, helping the model to converge to the more optimal solution.
@@ 5.5.2 Diffeomorphism Through Time
Table 4: Analysis of the effect of using discrete time steps on the Dice score and the percentage of negative Jacobian determinants.
@@
5.5.3 Effect of Discrete Time Steps
SGDIR can compute the deformation at an arbitrary time step using a continuous sampling from the time interval $[0,1]$. Here, we investigate the effect of using discrete time steps for training the model. The number of time steps indicates the number of intermediate time steps used during training. For instance, 0 time steps means that only the endpoint are used for registration (at time 1), and 8 time steps means that 8 intermediate time steps is used during training. $\infty$ is the continuous time sampling as done in our method. At the evaluation phase, we query the model at times 1 and -1 to warp the moving image towards the fixed image, and vice versa. Table 4 demonstrates that increasing the number of time samples in the interval $[0,1]$ improves the performance of the model both in terms of Dice score and $|det(J_{\phi})<0|$, while the best performance is achieved by continuous sampling. Figure 6 in Appendix 7.3 provides an illustrative inspection of this effect as well.

### v3

@@ 5 Experiments
5.1 Datasets and Metrics
To evaluate SGDIR and compare its performance with other methods, we consider four publicly available datasets on 3D MR brain scans widely used in the literature, namely OASIS [51], CANDI [52], LPBA40 [53], and IXI  [^1] https://brain-development.org/ixi-dataset/. To evaluate the performance of the model and for quantitative comparisons we use Dice score, 95th percentile Hausdorff Distance (HD95), and the percentage of negative Jacobian determinants. We provide the details of datasets, data preparations and evaluation metrics in Appendices 7.2 and 7.4.
@@ 5.2 Experimental Settings
In the OASIS dataset, out of 416 subjects, we use 256, 60, 100 subjects for training, validation, and testing, respectively. By random sampling we generate 1000 pairs for training, 100 pairs for validation and 1000 pairs for test. In the CANDI dataset, out of 117 subjects, we use 80, 11, 26 subjects for training, validation, and testing, respectively. We generate 400 pairs for training, 25 pairs for validation, and 300 pairs for test. For the LPBA40 dataset 190 pairs from 20 subjects are used as the training set, 10 pairs from 5 subjects are used as the validation set and 105 pairs from 15 subjects are used for the test set. Also, for the IXI dataset, out of 581 subjects, we use 450, 50, 81 subjects for training, validation, and testing, respectively, from which we create 1500, 100, 300 random pairs for training, validation, and test. We compare the performance of our model with a traditional optimization-based method, SyN [4], as one of the most successful methods for diffeomorphic image registration, learning-based methods including SYMNet [20], VoxelMorph [21], LapIRN [43], DiffuseMorph [27], PULPo [44], TransMorph-diff (TM-diff) [46], and optimization-based methods such as IDIR [8], NODEO [9], NePhi [39], DNVF and its variant Cas-DNVF [10]. We also compare our method with several non-diffeomorphic deformable methods including H-ViT [47], CorrMLP [54], CycleMorph [30], and TransMatch [55]. For SyN, we used the implementation in the DIPY package [56] and used cross-correlation as the loss function following the convention of the VoxelMorph. For other methods, except for DNVF/Cas-DNVF whose implementation was not available at the time of writing this paper, we used the codes available in their corresponding official repositories. The architecture of the UNet along with the time embedding modules is shown in Figure 1. We have provided our full implementation setup in Appendix 7.6.
@@ 5.3 Results
Tables 2, 2, 4, 4 provide a comprehensive comparison of the performance of SGDIR with other aforementioned methods on OASIS, CANDI, LPBA40, and IXI, respectively. These tables provide the performance of models in terms of Dice score, HD95 metric, and the percentage of negative Jacobian determinants. All the values are expressed as their average over the test set along with their standard deviation. For DNVF and Cas-DNVF due to unavailability of the implementation the results are directly taken from the corresponding paper. We consider the percentage of negative Jacobian determinants smaller than $10^{-5}\%$ negligible and show it as $\approx 0$ in the tables. Also, Figure 2 illustrates the qualitative comparison between our model and several other models on the OASIS dataset in terms of the warped image, warped segmentations, and the deformation grids. More visual results can be found in Appendix 7.9.
@@
Table 1: Comparison of the performance of SGDIR with other methods on the OASIS dataset. The values inside the parenthesis indicate the standard deviation.
Category Model Dice ($\uparrow$) HD95 ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 73.14 (1.32) 3.01 (0.28) 0.0531% (0.0241)
Learning VoxelMorph 75.93 (1.44) 2.31 (0.24) 0.0026% (0.0017)
SYMNet 79.39 (1.74) 1.93 (0.37) 0.0043% (0.0017)
LapIRN 78.80 (1.92) 1.87 (0.23) 0.0011% (0.0003)
DiffuseMorph 76.23 (2.23) 2.34 (0.44) 0.0551% (0.0131)
PULPo 77.36 (1.48) 2.05 (0.32) 0.0137% (0.0085)
CorrMLP 80.78 (1.33) 1.67 (0.33) 0.1162% (0.0310)
H-ViT 81.14 (1.31) 1.53 (0.26) 0.31% (0.012)
TransMatch 80.02 (1.14) 1.87 (0.34) 0.42% (0.093)
CycleMorph 77.32 (2.34) 1.97 (0.33) 0.22% (0.15)
TransMorph-diff 79.84 (1.93) 1.88 (0.23) 0.0012 (0.0006)
Optimization IDIR 72.31 (1.23) 2.75 (0.34) 0.0389% (0.0105)
NODEO 80.86 (1.14) 1.55 (0.12) 0.0105% (0.0041)
NePhi 77.89 (1.34) 1.77 (0.13) 0.0011% (0.0003)
DNVF 79.4 - 0.0015%
Cas-DNVF 81.5 - 0.0036%
Our Method SGDIR 82.23 (1.38) 1.44 (0.16) 0.0009% (0.0007)
@@
Table 2: Comparison of the performance of SGDIR with other methods on the CANDI dataset.
Category Model Dice ($\uparrow$) HD95 ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 73.60 (1.15) 2.13 (0.18) 0.0011% (0.0008)
Learning VoxelMorph 75.63 (1.78) 1.78 (0.24) 0.0011% (0.0009)
SYMNet 75.89 (1.13) 1.77 (0.24) 0.0015% (0.0008)
LapIRN 76.23 (1.76) 1.69 (0.21) 0.0004% (0.0002)
DiffuseMorph 75.63 (2.01) 1.83 (0.26) 0.0041% (0.0012)
PULPo 77.97 (1.43) 1.80 (0.17) 0.0097% (0.0011)
CorrMLP 78.88 (1.92) 1.78 (0.19) 0.0018% (0.0007)
H-ViT 80.63 (1.10) 1.66 (0.14) 0.0476% (0.0184)
CycleMorph 77.31 (2.09) 1.55 (0.13) 0.0024% (0.0011)
TransMorph-diff 77.45 (1.39) 1.51 (0.19) 0.0021% (0.0009)
Optimization IDIR 74.42 (1.12) 1.98 (0.11) 0.0113% (0.0081)
NODEO 78.02 (1.52) 1.76 (0.24) $\approx$ 0%
NePhi 78.89 (1.44) 1.73 (0.21) 0.0004% (0.0001)
DNVF 78.6 - 0.0003%
Our Method SGDIR 80.73 (1.25) 1.33 (0.20) $\mathbf{\approx 0}$%
Table 3: Comparison of the performance of SGDIR with other methods on the LPBA40 dataset.
Category Model Dice ($\uparrow$) HD95 ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 66.82 (3.17) 5.08 (0.34) 0.0194% (0.0072)
Learning VoxelMorph 70.72 (2.33) 4.02 (0.23) 0.0023% (0.0014)
SYMNet 69.98 (1.37) 4.06 (0.17) 0.0018% (0.0012)
LapIRN 71.88 (1.36) 3.98 (0.26) 0.0011% (0.0005)
DiffuseMorph 68.73 (2.88) 4.13 (0.32) 0.0032% (0.0011)
PULPo 68.34 (1.88) 4.36 (0.45) 0.0052% (0.0014)
CorrMLP 73.12 (1.06) 3.89 (0.21) 0.0315% (0.0021)
H-ViT 74.87 (1.73) 3.73 (0.19) 0.0891% (0.033)
TransMatch 71.52 (1.66) 4.01 (0.27) 0.0981% (0.0113)
CycleMorph 71.03 (1.54) 3.97 (0.32) 0.0413% (0.0087)
TransMorph-diff 72.91 (1.46) 3.83 (0.31) 0.0093% (0.0031)
Optimization IDIR 67.93 (2.03) 4.75 (0.43) 0.0013% (0.0008)
NODEO 72.31 (1.12) 3.96 (0.23) 0.0009% (0.0004)
NePhi 69.39 (1.98) 4.76 (0.74) 0.0103% (0.0045)
Our Method SGDIR 74.23 (2.15) 3.78 (0.13) 0.0007% (0.0003)
Table 4: Comparison of the performance of SGDIR with other methods on the IXI dataset.
Category Model Dice ($\uparrow$) HD95 ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 65.14 (2.09) 5.12 (0.66) 0.0201% (0.0096)
Learning VoxelMorph 70.76 (2.32) 4.62 (0.31) 0.0575% (0.0043)
SYMNet 71.21 (1.73) 4.46 (0.27) 0.0084% (0.0033)
LapIRN 72.71 (1.94) 4.24 (0.37) 0.0043% (0.0012)
DiffuseMorph 69.11 (2.14) 4.91 (0.76) 0.0175% (0.0053)
PULPo 69.24 (2.31) 4.83 (0.55) 0.0126% (0.0044)
CorrMLP 75.44 (1.32) 3.76 (0.31) 0.0192% (0.0064)
H-ViT 76.02 (1.45) 3.52 (0.24) 0.23% (0.07)
TransMatch 74.19 (1.24) 3.89 (0.66) 0.87% (0.026)
CycleMorph 70.39 (1.43) 4.73 (0.54) 0.47% (0.013)
TransMorph-diff 72.30 (2.01) 4.08 (0.43) 0.0049% (0.0004)
Optimization IDIR 68.25 (2.28) 5.01 (0.68) 0.0105% (0.0033)
NODEO 76.81 (1.09) 3.11 (0.14) 0.0013% (0.0008)
NePhi 73.81 (1.09) 4.14 (0.38) 0.0093% (0.0027)
Our Method SGDIR 77.98 (1.08) 3.02 (0.12) 0.0012% (0.0009)
Tables 2, 2, 4, 4 indicate that SGDIR outperforms competent diffeomorphic methods both in terms of Dice score, and $|det(J_{\phi})<0|$. SGDIR demonstrates 1.7%, 2.3%, 1.8%, and 1.5% improvements on Dice score with respect to the best evaluated diffeomorphic method over the OASIS, CANDI, LPBA40, and IXI datasets respectively. Among the non-diffeomorphic methods, H-ViT demonstrates the highest performance. The experiments demonstrate the SGDIR achieves competitive results against H-ViT in terms of Dice score while significantly improving the smoothness of deformation grid through $|det(J_{\phi})<0|$. This implies that the imposition of the semigroup property as a regularization in the time continuum of the deformation field can in fact lead to a precise and smooth deformation.
Among the compared methods, DiffuseMorph is also capable of generating a temporal deformation flow. Therefore, we have provided a visual comparison of warped images within 7 time steps for the OASIS dataset in Figure 3. In this example we have used the subject id 1 and 2 of the OASIS dataset as the fixed image and the moving image respectively. More visual results are provided in Appendix 7.9. Also, additional results on another dataset with a different modality is available in Appendix 7.8.
@@
Figure 3: The visual comparison of the continuous deformations of DiffuseMorph and SGDIR for OASIS dataset.
@@ 5.4 Ablation Study
Here we investigate the effect of some hyperparameters and design choices along with some consequent properties of SGDIR. More precisely, we investigate the effect of the semigroup regularization factor $\lambda$, the effect of continuous time scale against discrete time, and the topology-preserving property throughout the time interval. Since the behavior of the model is the same for all datasets, we provide the analysis only for the OASIS dataset.
@@ 5.4.1 Effect of Regularization Factor
Eq. 13 is responsible for imposing the semigroup property to the model and consequently, make the model produce topology-preserving non-folding deformations. Table 5 provides the effect of changing the factor $\lambda$ for the semigroup regularization on the Dice score and $|det(J_{\phi})<0|$. This table indicates that for a choice of a very large $\lambda$ we could achieve zero-folding deformations with a decrease in the dice score. Decreasing $\lambda$ gives more freedom to the model to increase the similarity of the deformed image and the fixed image in exchange for a larger folding percentage. An interesting property exhibited here is that drastically decreasing $\lambda$ (e.g., $\lambda=1$) is not helpful for achieving the highest Dice score. This indicates that the semigroup property is also helpful in narrowing down the trajectory space of the deformations, helping the model to converge to the more optimal solution.
@@ 5.4.2 Diffeomorphism Through Time
5.4.3 Effect of Discrete Time Steps
SGDIR can compute the deformation at an arbitrary time step using a continuous sampling from the time interval $[0,1]$. Here, we investigate the effect of using discrete time steps for training the model. The number of time steps indicates the number of intermediate time steps used during training. For instance, 0 time steps means that only the endpoint are used for registration (at time 1), and 8 time steps means that 8 equidistant intermediate time steps are used during training. $\infty$ is the continuous time sampling as done in our method. At the evaluation phase, we query the model at times 1 and -1 to warp the moving image towards the fixed image, and vice versa. Table 6 demonstrates that increasing the number of time samples in the interval $[0,1]$ improves the performance of the model both in terms of Dice score and $|det(J_{\phi})<0|$, while the best performance is achieved by continuous sampling. Figure 5 in Appendix 7.9 provides an illustrative inspection of this effect as well.
Table 6: Effect of the number time steps on the Dice score and the percentage of negative Jacobian determinants.
@@ 5.4.3 Effect of Discrete Time Steps
Table 7: Inference time comparison between SGDIR and other methods.
Model Inference Time (s) ()($\downarrow$)( ↓ )
SYMNet 0.41
VoxelMorph 0.41
TM-diff 0.39
PULPo 0.38
DiffuseMorph 5.77
H-ViT 0.31
NODEO 58.37
IDIR 149.24
NePhi 171.86
SGDIR 0.23
5.5 Computational Analysis
SGDIR is comprised of a rather simple architecture and avoids scaling and squaring integration. Here we have compared the inference time of SGDIR with other methods to evaluate how removing the integration can help the method to have faster inference time. The inference times are calculated for a single pair of brain MR scans. Table 7 compares the inference times. We note that among the learning-based diffeomorphic methods, SGDIR is 1.8 times as fast as SYMNet which employs scaling and squaring during the inference. DiffuseMorph requires us to take the iterative sampling steps of diffusion models and consequently, has much higher inference time. Due to existence of computationally expensive self- and cross-attentions, H-ViT demonstrates inferior inference time comparing to SGDIR. Finally, instance optimization-based methods which need to optimize the deformation grid for separate pairs, demonstrate orders of magnitude higher inference times, which is expected from such methods. More details on the number of parameters of SGDIR and its memory use can be found in Appendix 7.6 and in Table 9.